One morning, John drove 5 hours before stopping to eat. After lunch he increased his speed by 10 mph. If he completed a 430 mile trip in 8 hrs of driving time, how fast did he drive in the morning?
Answer by rajagopalan(158)

Let the morning speed =x
Driving time 5 hrs
Distance covered = speedxtime =5x........D1
Post lunch travel time =8-5=3
Post lunch speed=x+10
Post Lunch Distance covered = speedxtime =3(x+10)=3x+30......D2
Total distance covered=D1+D2=5x+3x+30
But 5x+3x+30 = 430
8x+30=430
8x=430-30
8x=400
x=400/8=50
The speed in the morning = 50 mph